The mimimum value of y=5x2 - 2x +1 is?? The answer is 4/5 but can you explain how?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Stdent,
Given function as:
y= 5x^2 – 2x + 1
Differntiate it once and equate to zero.
        So, dy/dx = 10x -2
         For minimum value
 
          10x-2=0
          Or, x= 1/5
         So, y= 5x1/25 – 2x1/5 + 1
                 = 1/5-2/5+ 1
                 = 4/5.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)